please answer all questions and show all working





2. Solve the the following differential equations (you answers will not have any arbitrary constants) (a) y+y - 2y =0 with
Answer #1

multiphy ear by -8-V3 2 and subtrald egn 2 from o - 5 – VB 5-V3 4+ (-5-3) 2 2 -5 tun -5-VB at С, со 2 2 _V3 Ci to =-5-813 2 G@ yll+Byl-gyo Acesallory ean is m² +8m -9--0 rootr are calculated or m²+gm-m-gro mimtg) -1 (mto) io mall -9 and im the generaegn @ yt yl-2y=0 auri sallang m² + m -2 o whose roots are given as - 1 tri ชา) 4X 1X (-2) 2x1 - 4 I Vito 2 < -) 13 2 mia (de☺ 67 To sulty zo 6m2-smt9=0 ausiltong ear in colore root are 51 Vare - 4x1x6 ma 2x6 s 25-24 12 sti 12 mi m흠 07P 모드봄 Ź and m2O yl +54 +3y=0 with y (0)4, 9 (0)2 o auseillany en is m?+sm+370 whose roots are -st/25-4x371 2X1 - 57/13 2 2 mia.-5+ VB m =